OpenStudy (anonymous):

Is the following system consistent or inconsistent? 8x + y + 4 = 0 y = -8x - 4

OpenStudy (hexagon001):

re arrange the first equation... if once you rearrange to have y on one side and the rest of the parts on the other..compare to the second equation.. if they look the same then they are consistent...

jimthompson5910 (jim_thompson5910):

Alternatively, plug in y = -8x - 4 into 8x + y + 4 = 0 to get 8x + y + 4 = 0 8x + (-8x - 4 + 4 = 0 8x - 8x - 4 + 4 = 0 I'm sure you see what the answer is from here

OpenStudy (anonymous):

So then it is consistent?

jimthompson5910 (jim_thompson5910):

yep, if you simplify fully, you'll get 0 = 0 That means you'll have infinitely many solutions, which makes the system consistent

OpenStudy (anonymous):

I really love how you guys explain everything on here. Thanks! :)

jimthompson5910 (jim_thompson5910):

yw